# Can LLMs be Good Graph Judge for Knowledge Graph Construction?

### v3

@@ 5.4 Generalization Capabilities of GraphJudge (RQ3)
To demonstrate the generalization abilities of GraphJudge, we conduct experiments in cross-dataset scenarios. We train it on GenWiki and then evaluate it on REBEL-Sub. As shown in Table 3, our method can still outperform baseline methods, which indicates GraphJudge has great capabilities of generalization across various corpus. This is because the ability to check facts with triple structure learned from graph judgement tasks can be generalized. It also suggests that GraphJudge once well trained on a general dataset, can be readily applied to diverse datasets with common knowledge.
@@ 7 Limitations
GraphJudge has the following limitations. First, even though we employ the LLMs act as both the extractor and the judge to improve the quality of constructed KGs, we still use the entity-level triples to construct KGs and there could be better knowledge units to form a better KG. Second, a more reasonable benchmark to evaluate the quality of constructed KGs should be proposed in the future. Currently, most of the work just utilize the ”ground truth” KGs to calculate the correctness and comprehensiveness of constructed KGs, However, the quality of ”ground truth” KGs may still deserve suspicion. So using a self-supervised approach to evaluate the KGs is in demands. We will research for more KG constructing and evaluating method to improve the performance of knowledge extraction.
@@ Appendix B Experimental Metrics
G-BLEU (G-BL): BLEU (Bilingual Evaluation Understudy)Papineni et al. (2002) is a metric for evaluating the quality of text which has been machine-translated from one natural language to another. Here we use this approach to determine the resemblance between the triple sentences in the ground-truth and predicted KGs, which is called G-BLEU. The formulas are shown in (16), (15), (12), (13), (14). is the maximum order of n-grams considered in the evaluation and we set , which is a default number in the Python package222https://pypi.org/project/bert-score/. is the brevity penalty, which is used to avoid giving too much credit to short translations. And we compute the accuracy, recall, and F1 score of each constructed KG against the ground-truth using G-BLEU, denoted as G-BL-Acc. G-BL-Recall, and G-BL-F1, respectively.
@@ Appendix D Graph Judgement Algorithm
| GenWiki-Hard | [Lower-Upper] | 53.33-100.0 | 68.00-100.0 |
@@
Appendix E Effect of ECTD Module
@@
Appendix F Knowledge Retention of ECTD Module and KG
@@
Appendix G Effect of KASFT Module
In this section, we extend the baseline models to explore their abilities to be a graph judge with the same experimental settings in Section 5.3. We extend baseline models to fine-tuned BERT, DeepSeek-V3, LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-70B, and GPT-4o. Compared with them, our proposed GraphJudge demonstrates consistent superiority in graph judgement tasks, which further proofs that the KASFT module can improve the capabilities of open-source LLMs as a graph judge. And neither fine-tuning a PLM with a smaller parameter size nor directly employing a powerful closed-source LLM can achieve a high accuracy on graph judgement tasks, which suggests the necessity to introduce our proposed GraphJudge.
@@ Appendix G Effect of KASFT Module
Appendix H Prompt Templates
As shown in Figure 8 and Figure 9, we demonstrate the prompt templates for the closed-source LLM to conduct relation extraction and for the open-source LLM to perform graph judgements on the results generated from the closed-source LLM. We also provide the prompt templates used to generate and answer MCQs in Appendix F.
Appendix I Case Study
In this section, we present an instance of constructing a KG from a document, achieved through the integration of a naive LLM (GPT-4o-mini) and our GraphJudge. We select a text-graph pair from the SCIERC dataset and contrast the results yielded by our approach with that of GPT-4o-mini. As shown in Table 7, the KG constructed by GPT-4o-mini with the given original document includes lots of meaningless triples. For example, <We, suggest, goal>, <We, suggest, evaluation criterion>, <We, present, measure>, <We, present, selection function>, etc. It is obvious that these triples do not convey any beneficial information that could be applied to subsequent tasks. And the triple <evaluation criterion, new, goal>does not even follow the general structure of triples, which means that the adjective word ”new” is generally not employed as a relational term within triples. The naive LLM have strong zero-shot ability to generate them but it does not have the capability to determine whether they are useful. However, there are no such triples in the KG constructed by our GraphJudge. On the one hand, this is because the triples without any useful information will be clarified as wrong triples by our fine-tuned LLM in graph judgement module. On the other hand, as demonstrated in the case, the document refined by ECTD module exhibits enhanced standardization and a reduction in irrelevant terms, for instance, terms such as ”-LRB-” and ”-RRB-” have been excluded as they are irrelevant to the document’s subject matter.

### v4

@@ 5.4 Generalization Capabilities of GraphJudge (RQ3)
To demonstrate the generalization abilities of GraphJudge, we conduct experiments in cross-dataset scenarios, which are training the LLM on GenWiki and then evaluate it on REBEL-Sub, training the LLM on REBEL-Sub and then evaluate it on GenWiki and SCIERC, respectively. As shown in Table 3, our method can still outperform baseline methods, which indicates GraphJudge has great capabilities of generalization across various corpus. This is because the ability to check facts with triple structure learned from graph judgement tasks can be generalized. It also suggests that GraphJudge once well trained on a general dataset, can be readily applied to diverse datasets with common knowledge. We also note that there is a gentle performance drop on SCIERC dataset, which is reasonable because there is less in-domain knowledge in the REBEL-Sub dataset than that in SCIERC dataset. And this result can further demonstrate the domain-specific knowledge within the training corpora is usefull in GraphJudge.
@@ 7 Limitations
GraphJudge has the following limitations. First, even though we employ the LLMs act as both the extractor and the judge to improve the quality of constructed KGs, we still use the entity-level triples to construct KGs and there could be better knowledge units to form a better KG. Second, a more reasonable benchmark to evaluate the quality of constructed KGs should be proposed in the future. Currently, most of the work just utilize the ‘ground truth’ KGs to calculate the correctness and comprehensiveness of constructed KGs, However, the quality of ‘ground truth’ KGs may still deserve suspicion. So using a self-supervised approach to evaluate the KGs is in demands. We will research for more KG constructing and evaluating method to improve the performance of knowledge extraction.
@@ Appendix A Datasets
Re-DocRED. Re-DocRED Tan et al. (2022) is a well annotated domain-specific dataset that is used in LKD-KGC Sun et al. (2025). It is a refined version of DocRED Yao et al. (2019), enhanced through reannotation of omitted relation triples. We follow the settings in LKD-KGC to use the Windows-centric subset containing domain knowledge about Windows Operation Systems. And because there is no enough samples in this dataset for us to train our model, we use the model trained on GenWiki to evaluate the performance of GraphJudge on this dataset.
@@ Appendix B Experimental Metrics
G-BLEU (G-BL): BLEU (Bilingual Evaluation Understudy)Papineni et al. (2002) is a metric for evaluating the quality of text which has been machine-translated from one natural language to another. Here we use this approach to determine the resemblance between the triple sentences in the ground-truth and predicted KGs, which is called G-BLEU. The formulas are shown in (16), (15), (12), (13), (14). is the maximum order of n-grams considered in the evaluation and we set , which is a default number in the Python package222https://pypi.org/project/bert-score/. BP is the brevity penalty, which is used to avoid giving too much credit to short translations. And we compute the accuracy, recall, and F1 score of each constructed KG against the ground-truth using G-BLEU, denoted as G-BL-Acc. G-BL-Recall, and G-BL-F1, respectively.
@@ Appendix D Graph Judgement Algorithm
| GenWiki | [Lower-Upper] | 53.33-100.0 | 68.00-100.0 |
@@
Appendix E Entity Coverage of the LLM Extraction
| Dataset | Method | Recall | Acc |
| REBEL-Sub | w/o GJ | 0.9784 | 0.3709 |
| w/ GJ | 0.9411 | 0.5495 | |
| GenWiki-Hard | w/o GJ | 0.9365 | 0.7932 |
| w/ GJ | 0.9018 | 0.8343 | |
| SCIERC | w/o GJ | 0.9392 | 0.6916 |
| w/ GJ | 0.9187 | 0.7176 |
To verify the comprehensiveness of extracted entities by the LLM, we evaluate the entity coverage as well as accuracy across three datasets. In detail, we use the semantic metric of Recall and Accuracy based on BertScore to measure the entity coverage and accuracy, similar to the metric G-BS described in the paper. As shown in Table 7, we report the coverage as well as accuracy of the extracted entities before (w/o GJ) and after (w/ GJ) the GJ module. And here are the insights: (1) The entities extracted from the original text have high coverage and low accuracy, which verifies that the extracted entities can provide good foundation for both denoising and relation extraction. (2) The entities within the triples filtered by GJ module still keep a high coverage and have a significantly higher accuracy. It further validates the effectiveness of our GJ module.
Appendix F Effect of ECTD Module
@@
Appendix G Knowledge Retention of ECTD Module and KG
@@
Appendix H Effect of KASFT Module
In this section, we extend the baseline models to explore their abilities to be a graph judge with the same experimental settings in Section 5.3. As shown in Figure 7, We extend baseline models to fine-tuned BERT, DeepSeek-V3, LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-70B, and GPT-4o. Compared with them, our proposed GraphJudge demonstrates consistent superiority in graph judgement tasks, which further proofs that the KASFT module can improve the capabilities of open-source LLMs as a graph judge. And neither fine-tuning a PLM with a smaller parameter size nor directly employing a powerful closed-source LLM can achieve a high accuracy on graph judgement tasks, which suggests the necessity to introduce our proposed GraphJudge.
Appendix I Prompt Templates
As shown in Figure 8 and Figure 11, we demonstrate the prompt templates for the closed-source LLM to conduct relation extraction and for the open-source LLM to perform graph judgements on the results generated from the closed-source LLM. We also provide the prompt templates used to generate and answer MCQs in Appendix G.
Appendix J Case Study
In this section, we present an instance of constructing a KG from a document, achieved through the integration of a naive LLM (GPT-4o-mini) and our GraphJudge. We select a text-graph pair from the SCIERC dataset and contrast the results yielded by our approach with that of GPT-4o-mini. As shown in Table 8, the KG constructed by GPT-4o-mini with the given original document includes lots of meaningless triples. For example, <We, suggest, goal>, <We, suggest, evaluation criterion>, <We, present, measure>, <We, present, selection function>, etc. It is obvious that these triples do not convey any beneficial information that could be applied to subsequent tasks. And the triple <evaluation criterion, new, goal>does not even follow the general structure of triples, which means that the adjective word ‘new’ is generally not employed as a relational term within triples. The naive LLM have strong zero-shot ability to generate them but it does not have the capability to determine whether they are useful. However, there are no such triples in the KG constructed by our GraphJudge. On the one hand, this is because the triples without any useful information will be clarified as wrong triples by our fine-tuned LLM in graph judgement module. On the other hand, as demonstrated in the case, the document refined by ECTD module exhibits enhanced standardization and a reduction in irrelevant terms, for instance, terms such as ‘-LRB-’ and ‘-RRB-’ have been excluded as they are irrelevant to the document’s subject matter.